OpenStudy (anonymous):

Check my answer? y=-x^10 is a even monomial function. True or False? I say True because the exponent is even. Am I right?

OpenStudy (accessdenied):

Even means that y for x is the same as y for -x. That checks out because (-x)^10 = x^10. It is a monomial because it only has one term as well. So, true sounds good. :)

OpenStudy (anonymous):

So for the y for x = y for -x, you're saying that it's even because even though the coefficient is negative, the coefficient would become positive after being multiplied by itself an even amount of times? Please tell me if that made no sense to you. I think my wording was a bit confusing.

OpenStudy (accessdenied):

The coefficient is not affected, actually. I just omitted it for brevity, but: \( y (x) = -x^{10} \) \( y (-x) = -(-x)^{10} = - x^{10} = y(x) \) This is the definition of an even function: \( y(x) = y(-x) \)

OpenStudy (anonymous):

Okay I get it. Thank you again. :)

OpenStudy (accessdenied):

You're welcome!